OpenStudy (anonymous):

If a piece of sea floor has moved 50 km in 5 million years, what is the yearly

OpenStudy (aum):

50 km in 5 million years = \[ \frac{50}{5\times 10^6} \frac{\text{Km}}{\text{year}} = \frac{50\times10^5}{5\times 10^6} \frac{\text{cm}}{\text{year}} = 1\frac{\text{cm}}{\text{year}} \]